# CoBRA: A Universal Strategyproof Confirmation Protocol for Quorum-based Proof-of-Stake Blockchains

### v2

@@ Abstract
We present a formal analysis of quorum-based State Machine Replication (SMR) protocols in Proof-of-Stake (PoS) payment systems under a hybrid threat model comprising honest, Byzantine, and rational validators. Our analysis of traditional quorum-based protocols establishes two fundamental impossibility results: (1) in partially synchronous networks, no quorum-based protocol can achieve SMR when rational and Byzantine validators comprise more than $1/3$ of participants, and (2) even in synchronous networks, SMR remains impossible when rational and Byzantine validators comprise $2/3$ or more of participants.
@@
Given these limitations, we propose two complementary solutions that achieve SMR in the presence of less than $2/3$ Byzantine and rational validators. To achieve this, we first introduce a protocol that enforces an upper bound on the total volume of transactions finalized within any time window $\Delta$ and prove that this bound is necessary for secure SMR protocols in our model. Second, we present the strongest chain rule, which enables efficient finalization of transactions when the majority of honest participants provably support the SMR execution. Through empirical analysis of Ethereum and Cosmos networks, we demonstrate that validator participation for over $99\%$ of blocks exceeds the required $5/6$ threshold, establishing the practical feasibility of our solution in production PoS systems.
1 Introduction
State Machine Replication (SMR) protocols are the backbone of fault-tolerant distributed systems, ensuring agreement on an ordered sequence of operations among participants. Traditionally, these protocols classify participants as either honest (strictly following the protocol), or Byzantine (deviating arbitrarily). However, this binary model fails to capture the behavior of rational participants, who strategically deviate only when it maximizes their utility.
This distinction is particularly critical in Proof-of-Stake (PoS) blockchains, which collectively secure $0.5$ trillion dollars. Participation in these systems is open and financially motivated, with validators staking assets to partake in consensus and earn rewards. While this incentive structure is designed to encourage honest behavior, it also introduces new attack vectors: profit-maximizing validators may deviate to increase their gains. Worse, external financial markets, such as derivative markets, may create additional incentives for validators to undermine the system. Capturing this multi-dimensional behavior requires a more nuanced model that accounts for honest, Byzantine, and rational validators. Ensuring the economic security of quorum-based SMR protocols (e.g., PBFT [1], HotStuff [2], Tendermint [3]) within this extended model is essential to safeguard existing PoS blockchains.
Several recent works examine the economic security of quorum-based PoS blockchains, e.g., [4, 5, 6, 7, 8, 9]. However, they either, like [6] adopt a different extended model and rely on a random beacon, which provides only probabilistic security. Or, in contrast, similar to [4] consider only Byzantine and honest validators but solely focus on slashing misbehaving validators while protecting honest stake, without accounting for liquid token transfers, which adversaries can exploit. Or, like [7], primarily focus on achieving economic safety after recovering from attacks. Yet, none of these works address the following fundamental question: Can existing PoS quorum-based SMR protocols be adapted to provably withstand both Byzantine and rational validators?
1.1 Contributions
We answer this question in the affirmative and introduce the first transformation that enables optimal rational resilience in PoS blockchains. We make the following contributions:
  1. 1.
    Impossibility Results: We prove that no quorum-based SMR protocol can be resilient when Byzantine and rational validators together exceed critical thresholds:
    • Partial Synchrony: Quorum-based SMR is impossible when Byzantine and rational validators collectively control $1/3$ or more of the validator set.
    • Synchrony: Quorum-based SMR is impossible when Byzantine and rational validators collectively control more than $2/3$ of the validator set.
  2. 2.
    Protocol Transformation: We introduce CoBRA (Correct over Byzantine Rational Adversary), a transformation that enhances existing quorum-based PoS protocols to tolerate rational behavior . Our approach comprises:
    • A finality gadget that detects equivocations and prevents rational validators from benefiting from forking. Our finality gadget safeguards honest validators as well as clients that do not actively monitor the SMR protocol execution.
    • A stake-bounded finalization rule that enables finalizing a total transaction volume equal to the stake validators lock in the system before a synchronous bound elapses.
    • A strongest-chain finalization rule that enables unbounded transaction throughput when at least $5/6$ of the validators actively follow the protocol.
    • An incentive mechanism consisting of (i) slashing conditions to disincentivize rational validators to equivocate in collusion with the Byzantine ones; (ii) a novel reward mechanism to incentivize rational validators to actively participate in the SMR protocol, and not engage in censoring or silencing honest proposers.
  3. 3.
    Incentive-based Security Analysis: We prove that CoBRA can securely transform the class of quorum-based SMR protocols to withstand rational participants, effectively reducing the required fraction of honest validators while maintaining security guarantees.
  4. 4.
    Empirical Validation: We evaluate validator participation rates in Ethereum and Cosmos, revealing they exceed the strongest-chain threshold for over 99% of the blocks, highlighting CoBRA is well-suited for deployment in existing systems.
2 Preliminaries
2.1 Model and Assumptions
@@
Proof-of-Stake-based blockchain
A PoS blockchain implements a distributed ledger protocol [10] (i.e., a state machine replication protocol), where participants are selected to create and validate new blocks based on the amount of coins (cryptocurrency) they hold and lock up as stake. The blockchain serves as a payment system in which participants hold and transfer fungible assets, referred to as coins. We also discuss how to extend this model to general-purpose smart contract platforms in Section 7. Each party is associated with a private-public key pair which forms its client identity in the system. The parties that actively participate in maintaining the state of the system (i.e., run the SMR) are known as validators. To do so, they lock coins (i.e., stake) and process transactions, providing clients an ordered sequence of confirmed transactions, referred to as the transaction ledger.
We assume the existence of a genesis block $G$, which contains the initial stake distribution and the identities of the validators participating in the SMR protocol. For the remainder of this paper, we assume a static validator setting, where the stake distribution is known in advance to all participants. In Section 7 we discuss how we can extend our system to a dynamic validator setting.
@@
Unlike validators, the number of active clients is not known to all parties. Clients have public keys, which they can use to issue authenticated transactions. We assume clients are silent [11], meaning they neither actively interact with nor observe the execution of the SMR protocol among validators. Instead, clients query a subset of validators to obtain confirmation of any transaction (i.e., the transaction ledger) from their responses.
@@
Time progresses in discrete logical units indexed as $u=0,1,2,\dots$. We assume that the maximum clock offset between any two parties is bounded, as any such offset can be absorbed into the network delay bound.
@@
We assume that all participants are connected through perfect point-to-point communication channels. We consider two communication models: partial synchrony and synchrony. In the former, we show impossibility results; in the latter, we present both impossibility and our main results.
@@
When applying our transformation protocol to partially synchronous SMR protocols, we provide an alternative interpretation of our synchronous network model in Section 7, following prior work [12, 8].
@@
We consider a mixed model where validators are honest, Byzantine, or rational. Honest validators always follow the protocol. Byzantine validators deviate arbitrarily from the protocol. We assume a Byzantine adversary controls at most $f=(n-1)/3$ validators, and does not have any liquid coins, i.e., all their coins are staked to be able to maximize their influence.
Rational validators, on the other hand, will deviate from the protocol specification if and only if doing so maximizes their utility. We define a party’s utility as its monetary profit, calculated as the rewards from participation minus any potential penalties (e.g., slashing). For simplicity, we assume no loss in opportunity cost. We denote the number of rational validators in the system by $k$.
@@
We make the following assumptions on rational validators: (a) The cost of operating a validator is assumed to be negligible. Consequently, any positive reward provides sufficient incentive for rational validators to participate. This assumption simplifies the analysis by canceling out the constant operational costs of running a PoS validator (e.g., bandwidth, CPU, storage) with a corresponding minimum constant reward. (b) Second, the cost of voting for a block proposed by another validator (e.g., adding a signature) is negligible and does not decrease their utility. (c) Third, a rational validator deviates from the protocol only if such deviation results in strictly greater utility, as devising a strategy to deviate may incur additional costs to running the default client. For example, a rational validator will censor or double spend only if they can secure a bribe larger than their potential loss due to slashing. This stands in contrast to Byzantine validators, who may deviate arbitrarily, capturing scenarios that fall outside the scope of our utility-based PoS model. (d) A rational validator considers a strategy only if the resulting utility is guaranteed within the transaction ledger, as we do not account for external trustless mechanisms  [^1] Any deviation from this assumption, such as validators influenced by external trustless mechanisms is modeled as Byzantine behavior.. Promises or unverifiable commitments from the adversary regarding future actions are not considered credible and, thus, will not influence the validator’s strategy. Given the appropriate incentives, however, rational validators may collude with other rational and Byzantine validators.
2.2 Definitions and Properties
@@
At a high level, in an SMR protocol, validators must perform two primary tasks: order a set of transactions [^2] Validators execute the transactions in the agreed order to acquire the state of the state machine. and reply to clients. Traditional SMR protocols such as PBFT [1] and HotStuff [2] assume an honest supermajority of validators – specifically $12f+$ 1 out of the total $13f+$ 1 validators must be honest. Under this assumption, clients can extract the confirmation of transactions simply by waiting for responses from $f+1$ validators. In contrast, our setting introduces rational validators and assumes only an honest minority. For this reason, we separate these two responsibilities: (i) ordering transactions and (ii) replying to clients.
@@
Given as input transactions, the validators execute an SMR protocol $\Pi$ and output an ordered sequence of confirmed transactions, known as the transaction ledger. We say that $\Pi$ solves the SMR problem when: i) correct validators agree in a total order [13], ii) the transaction ledger is certifiable.
@@
  • Liveness. Any valid transaction witnessed by all correct validators, will eventually be included in the transaction ledger of every correct node.
@@
Traditional SMR protocols [14, 1, 15] are secure when considering $f$ Byzantine validators and $n-f$ correct validators. We say that a protocol satisfying this condition is $(n,f)$-resilient.
@@
A protocol $\Pi$ executed by $n$ validators is $(n,f)$-resilient if it remains secure in the presence of up to $f$ Byzantine validators, assuming that the remaining $n-f$ validators follow the protocol correctly.
@@
A protocol $\Pi$ executed by $n$ validators is $(n,k,f)$-resilient if it remains secure in the presence of $f$ Byzantine and $k$ rational validators, assuming that the remaining $n-f-k$ validators follow the protocol correctly.
@@
We focus on a family of SMR protocols that progress by collecting certificates of $q$ votes across multiple phases, similar to several BFT SMR protocols (e.g. [1, 14]). We define a quorum certificate (QC) as a certificate consisting of $q$ votes. A protocol is q-commitable if a QC on a specific message $m$ (e.g., $m=\texttt{commit}$ in PBFT [1]) is both a necessary and sufficient condition for validators to commit a transaction (or block) $b$ to their local ledger.
@@
We study SMR protocols in the context of proof-of-stake (PoS) blockchains, where validators batch transactions into blocks, as in HotStuff [2]. Our objective is to determine when an $(n,f)$-resilient SMR protocol can be transformed into an $(n,k,f)$-resilient PoS blockchain. Specifically, we present a transformation that converts any $(n,f)$-resilient $q$-certifiable PoS blockchain [1, 14, 15] into an $(n,k,f)$-resilient PoS blockchain. At the core of this transformation, we introduce a new certification mechanism for finalizing transactions, ensuring security under our hybrid model.
@@
2.3 SMR game
In this section, we relate game-theoretic rational behavior to protocol-level resilience guarantees. While $(n,k,f)$-resilience specifies that a protocol remains safe and live when at most $f$ validators are Byzantine and $k$ are rational, it does not explain why rational validators would not prefer to violate the protocol properties. To address this, we model the protocol execution as a finite game and introduce the notion of a rationally robust protocol. This notion ensures that rational validators have no incentive to deviate in ways that compromise safety or liveness, regardless of the strategies thereby ensuring the protocol’s resilience.

### v3

@@ Abstract.
The integrity of many Proof-of-Stake (PoS) payment systems relies on quorum-based State Machine Replication (SMR) protocols mitigating diverse adversarial attacks. While traditional analyses assume a purely Byzantine threat model, practical security depends on resilience against both arbitrary faults and strategic, profit-driven actors. In this work, we study this challenge by analyzing SMR protocols under a hybrid threat model comprising honest, Byzantine, and rational validators. We first establish the fundamental limitations of traditional consensus mechanisms, proving two impossibility results: (1) in partially synchronous networks, no quorum-based protocol can achieve SMR when rational and Byzantine validators collectively exceed $1/3$ of the participants; and (2) even under synchronous network assumptions, SMR remains unattainable if this coalition comprises more than $2/3$ of the validator set.
@@
Motivated by these boundaries, we develop a protocol that achieves SMR in the presence of up to $1/3$ Byzantine and $1/3$ rational validators. Our approach relies on two complementary mechanisms to circumvent our impossibility results. First, we introduce a protocol constraint that enforces an upper bound on the transaction volume finalized within any time window $\Delta$, and we prove that such a bound is a necessary condition for security. Second, we define the strongest chain rule, a finality gadget that enables instant transaction confirmation when a supermajority of participants provably supports the execution. We validate the feasibility of this design through an empirical analysis of the Ethereum and Cosmos networks, demonstrating that validator participation exceeds the required $5/6$ threshold in over $99\%$ of observed blocks.
Finally, we explore adversarial settings with $\geq 1/3$ Byzantine validators. There we provide a recovery mechanism capable of restoring safety and liveness following consistency violations, even when Byzantine validators control up to $5/9$ of the stake and rational actors control $1/9$. This mechanism ensures security by guaranteeing that any clients incurring provable economic losses due to safety faults are fully reimbursed.
Consensus, Blockchain, Proof-of-Stake, Accountability, Recovery
conference: ; $c$s: Theory of computation Algorithmic game theory and mechanism designccs: Computer systems organization Dependable and fault-tolerant systems and networksccs: Security and privacy Distributed systems security
1. Introduction
State Machine Replication (SMR) protocols are the backbone of secure distributed ledgers, ensuring agreement on an ordered sequence of transactions among distributed participants. Traditional SMR protocols model participants as either honest, strictly following the protocol, or Byzantine, deviating arbitrarily. Quorum-based Byzantine Fault Tolerant (BFT) protocols, such as PBFT (Castro and Liskov, 2002), HotStuff (Yin et al., 2018), and Tendermint (Buchman et al., 2018), operate under partial synchrony and guarantee safety and liveness provided that less than $1/3$ of the participants are Byzantine. Once this threshold is exceeded, safety violations are unavoidable.
While the Byzantine model is deliberately pessimistic, the complementary assumption that a $2/3$ supermajority of participants behaves honestly at all times is particularly strong. This is evident in modern Proof-of-Stake (PoS) blockchains, where participation is economically motivated. Validators stake assets to earn rewards from block production, transaction fees, and external financial positions, collectively securing hundreds of billions of dollars in value. In this setting, participants are not altruistic entities but profit-maximizing agents. Deviations from the protocol may therefore be strategic rather than arbitrary, arising whenever doing so increases expected utility. As a result, violations of the honest supermajority assumption may stem not only from malicious behavior, but also from economically rational behavior.
This observation raises a fundamental question for the design of secure distributed ledgers: What assurances can we provide when fewer than $2/3$ of the participants behave honestly? In particular, can we design protocols that both (i) incentivize rational, profit-maximizing validators to follow the protocol, and (ii) recover from safety violations without imposing irreversible economic losses on clients, even when more than $1/3$ of the participants behave Byzantine?
1.1. Related Work
A substantial body of work has studied consensus under violations of the classic $2/3$ honest supermajority assumption. Broadly, this literature falls into two main categories.
Rational and hybrid fault models.
A first line of work studies consensus under hybrid fault models distinguishing honest, rational, and Byzantine participants. BAR (Aiyer et al., 2005) formalizes this setting, and subsequent works extend it to incentive-compatible and equilibrium-based consensus, ensuring honest behavior is a Nash equilibrium under appropriate rewards and penalties (Amoussou-Guenou et al., 2019; McMenamin et al., 2021; Lev-Ari et al., 2019). These approaches typically assume rational participants act independently, ignoring collusion. This assumption is problematic in open, financially motivated environments such as Proof-of-Stake blockchains, validators often coordinate in pools, making deviations like double-spending profitable when executed jointly.
To capture collusions, Abraham et al. (Abraham et al., 2006, 2008) formalize solution concepts for strategic interaction between rational and Byzantine participants in secret sharing and multi-party computation. Building on this, TRAP (Ranchal-Pedrosa and Gramoli, 2022) introduces a baiting game to hold misbehaving nodes accountable, solving one-shot consensus in a partially synchronous model for $n>\max\left(\frac{3}{2}k+3f,2(k+f)\right)$ where $k$ in the number of rational and $f$ the number of Byzantine participants. While one-shot consensus allows bounding the maximum double-spend, this does not extend to multi-shot SMR, where repeated strategic behavior can accumulate unbounded economic damage.
Accountability and recovery.
A second line of work studies consensus protocols that tolerate violations of the classic $n/3$ Byzantine threshold by weakening safety guarantees and providing accountability (Sheng et al., 2021; Civit et al., 2021, 2023) and recovery mechanisms (Lewis-Pye and Roughgarden, 2025; Ranchal-Pedrosa and Gramoli, 2024; Gong et al., 2025).
Accountable protocols focus on detecting equivocation and identifying misbehaving validators, but do not address recovery. As such, accountability alone does not restore safety or protect clients from economic losses incurred during safety violations.
Recoverable consensus protocols address this limitation by incorporating a mechanism that enables the system to internally restore a consistent state after safety violations, even when more than $1/3$ of the participants behave maliciously (Gong et al., 2025; Lewis-Pye and Roughgarden, 2025; Ranchal-Pedrosa and Gramoli, 2024; Sridhar et al., 2023b). However, these protocols either assume communication between clients to detect conflicting certificates (Sridhar et al., 2023b) or provide no economic guarantees to clients (Gong et al., 2025; Lewis-Pye and Roughgarden, 2025; Ranchal-Pedrosa and Gramoli, 2024): transactions accepted during safety violations may be reverted without compensation.
As a result, existing work does not resolve our fundamental question. There is no SMR protocol that (i) incentivizes profit-maximizing participants to follow the protocol and (ii) provides economic security to clients when the classic $1/3$ Byzantine threshold is exceeded.
1.2. Our Contributions
This paper presents CoBRA, which closes this gap. It is the first generic transformation for quorum-based SMR that provides optimal rational resilience when the Byzantine parties are less than 1/3 and economic protection for clients when they are more than 1/3 and less than 5/9. Table 1 compares CoBRA to prior work, highlighting the trade-offs in fault tolerance, recovery guarantees, and client-level economic protection.
Impossibility results.
We establish fundamental limits for quorum-based SMR under a hybrid fault model with honest, Byzantine, and rational validators, and non-interactive clients:
  • Under partial synchrony, we show that quorum-based SMR is impossible once Byzantine and rational validators together control at least $1/3$ of the validator set. This bound is tight and generalizes the classic Byzantine threshold to economically motivated deviations.
  • Under synchrony, we further prove that SMR remains impossible when this coalition exceeds $2/3$.
Protocol transformation.
We introduce CoBRA, a generic transformation applicable to existing quorum-based PoS protocols, yielding three operational regimes:
  • Standard regime: In partial synchrony, CoBRA preserves the safety and liveness guarantees of the underlying SMR protocol against up to $1/3$ Byzantine validators.
  • Rational-resilient regime: Assuming a (possibly large) known synchronous bound $\Delta^{*}$ on message delays, CoBRA ensures safety and liveness even when, in addition to $1/3$ Byzantine validators, up to another $1/3$ of the participants are rational and may engage in strategic collusion. This is achieved by modifying only the finalization rule of the underlying protocol.
  • Recoverable economic-security regime: Assuming a synchronous bound $\Delta^{*}$ when Byzantine behavior exceeds the standard $1/3$ threshold, CoBRA enables recovery from consistency violations and guarantees full client reimbursement. CoBRA provides these guarantees without requiring minting new coins or client participation in the recovery procedure.
2. Preliminaries
2.1. Model and Assumptions
@@
Proof-of-stake blockchain
A PoS blockchain implements a distributed ledger protocol (Garay et al., 2024) (i.e., a state machine replication protocol), where participants are selected to create and validate new blocks based on the amount of coins (cryptocurrency) they hold and lock up as stake. The blockchain serves as a payment system in which participants hold and transfer fungible assets, referred to as coins. We also discuss how to extend this model to general-purpose smart contract platforms in Section 8. Each party is associated with a private-public key pair which forms its client identity in the system. The parties that actively participate in maintaining the state of the system (i.e., run the SMR) are known as validators. To do so, they lock coins (i.e., stake) and process transactions, providing clients an ordered sequence of confirmed transactions, referred to as the transaction ledger.
We assume the existence of a genesis block $G$, which contains the initial stake distribution and the identities of the validators participating in the SMR protocol. For the remainder of this paper, we assume a static validator setting, where the stake distribution is known in advance to all participants. In Section 8, we discuss how we can extend our system to a dynamic validator setting.
@@
Unlike validators, the number of active clients is not known to all parties. Clients have public keys, which they can use to issue authenticated transactions. We assume clients are silent (Sridhar et al., 2024), meaning they neither actively interact with nor observe the execution of the SMR protocol among validators. Instead, clients query a subset of validators to obtain confirmation of any transaction (i.e., the transaction ledger) from their responses.
@@
Time progresses in discrete logical units indexed as $u=0,1,2,\dots$. Each participant has their own local clock. We assume that the maximum clock offset between any two parties is bounded, as any such offset can be absorbed into the network delay bound.
@@
We assume that all participants are connected through perfect point-to-point communication channels. We consider two communication models: partial synchrony and synchrony. In the former, we show impossibility results; in the latter, we present both an impossibility and our main results.
@@
We consider a mixed model where validators are honest, Byzantine, or rational. Honest (or correct) validators always follow the protocol. Byzantine validators may deviate arbitrarily from the protocol. We assume a Byzantine adversary does not have any liquid coins, i.e., all their coins are staked to be able to maximize their influence. Rational validators, on the other hand, will deviate from the protocol specification if and only if doing so maximizes their utility. We define a party’s utility as its monetary profit, calculated as the rewards from participation minus potential penalties (e.g., slashing).
We denote the number of Byzantine and rational validators by $f^{*}$ and $k$, respectively. In Section 3, we derive the feasible bounds for $f^{*}$ and $k$. In our solution, we consider $n=3f+1$, where $f$ represents the maximum number of Byzantine validators tolerated by standard BFT protocols, e.g., (Abraham et al., 2018; Buchman et al., 2018; Castro and Liskov, 2002). In Section 4, we specify that $f^{*}\leq f$ and $k\leq 2f-f^{*}$. Then, in Section 5, we extend our solution to Byzantine failures beyond the classic threshold, in particular, $f+1\leq f^{*}<5n/9$ and $k<2n/3-f^{*}$.
@@
We make the following assumptions on rational validators: (a) We assume that participation rewards exceed both the opportunity cost and the cost of operating a validator. Thus, these costs are treated as negligible, simplifying the analysis by offsetting the constant operational expenses (e.g., bandwidth, CPU, storage) or voting for a block proposed by another validator (e.g., adding a signature) and the constant opportunity cost with a minimum constant reward. In practice, rational validators account for these costs beforehand and conclude that participation is beneficial. (b) A rational validator considers a strategy only if the resulting utility is guaranteed within the transaction ledger, as we do not account for external trustless mechanisms. Promises or unverifiable commitments from the adversary regarding future actions are not considered credible and, thus, will not influence the validator’s strategy. Given the appropriate incentives, however, rational validators may collude with other rational and Byzantine validators. (c) A rational validator deviates from the protocol only if such deviation results in strictly greater utility, as devising a strategy to deviate may incur additional costs to running the default client. For example, a rational validator will censor or double spend only if they can secure a bribe larger than their potential loss due to slashing.
2.2. Definitions and Properties
@@
At a high level, in an SMR protocol, validators must perform two primary tasks: order a set of transactions which they execute to extract the state, and reply to clients. Traditional SMR protocols such as PBFT (Castro and Liskov, 2002) and HotStuff (Yin et al., 2018) assume an honest supermajority of validators – specifically $2f+1$ out of the total $3f+1$ validators must be honest. Under this assumption, clients can extract the confirmation of transactions simply by waiting for responses from $f+1$ validators. In contrast, our setting introduces rational validators and assumes only an honest minority. For this reason, we separate these two responsibilities: (i) ordering transactions, and (ii) replying to clients.
@@
Given as input transactions submitted by clients, the validators execute an SMR protocol $\Pi$ and output an ordered sequence of confirmed transactions, known as the transaction ledger. We say that $\Pi$ solves the SMR problem when: (i) correct validators agree in a total order (Défago et al., 2004), (ii) the transaction ledger is certifiable to clients.
@@
  • Liveness. Any valid transaction witnessed by all correct validators, will eventually be included in the transaction ledger of every correct validator.
@@
We emphasize that total order and certifiability are distinct guarantees. Some protocols achieve total order but not certifiability (e.g., Dolev–Strong (Dolev and Strong, 1983) for $f>n/2$ Byzantine faults). To satisfy both properties, validators and clients may adopt different confirmation rules, in line with the flexible BFT paradigm (Malkhi et al., 2019).
@@
Traditional SMR protocols (Abraham et al., 2018; Castro and Liskov, 2002; Gelashvili et al., 2022) are secure when considering $f$ Byzantine validators and nfn-f correct validators. We say that a protocol satisfying this condition is $(n,f)$-resilient.
@@
A protocol $\Pi$ executed by $n$ validators is $(n,f)$-resilient if it remains secure in the presence of up to $f$ Byzantine validators, assuming that the remaining nfn-f validators follow the protocol correctly.
@@
A protocol $\Pi$ executed by $n$ validators is $(n,k,f)$-resilient if it remains secure in the presence of $f$ Byzantine and $k$ rational validators, assuming that the remaining nfkn-f-k validators follow the protocol correctly.
@@
We focus on a family of SMR protocols that progress by collecting certificates of $q$ votes across multiple phases, similar to several BFT SMR protocols (e.g. (Castro and Liskov, 2002; Abraham et al., 2018)). We define a quorum certificate (QC) as a certificate consisting of $q$ votes. A protocol is q-commitable if a QC on a specific message $m$ (e.g., $m=\texttt{commit}$ in PBFT (Castro and Liskov, 2002)) is both a necessary and sufficient condition for validators to commit a transaction (or block) $b$ to their local ledger.
@@
We study SMR protocols in the context of proof-of-stake (PoS) blockchains, where validators batch transactions into blocks, as in HotStuff (Yin et al., 2018). Our objective is to determine when an $(n,f)$-resilient SMR protocol can be transformed into an $(n,k,f)$-resilient PoS blockchain. Specifically, we present a transformation that converts any $(n,f)$-resilient $q$-certifiable PoS blockchain (Castro and Liskov, 2002; Abraham et al., 2018; Gelashvili et al., 2022) into an $(n,k,f)$-resilient PoS blockchain. At the core of this transformation, we introduce a new certification mechanism for finalizing transactions, ensuring security under our hybrid model.
@@
Recovery and economic restitution
A protocol that is $(n,k,f)$-resilient constitutes a complete solution to SMR: when at most $f$ validators are Byzantine and at most $k$ are rational, the protocol preserves both safety and liveness. We now consider executions in which the actual number of Byzantine validators $f^{*}$ exceeds the resilience threshold $f$ of an $(n,f)$-resilient system. In such executions, the protocol may suffer both safety and liveness violations.
Liveness attacks cannot, in general, be deterred, since Byzantine validators may simply refuse to participate. For this reason, the literature on recoverable consensus typically assumes an adversary that is willing to participate in the protocol (e.g., to earn participation rewards) but may opportunistically violate safety; this type of faults are called a-b-c Byzantine (Malkhi et al., 2019; Ranchal-Pedrosa and Gramoli, 2024; Gong et al., 2025).
In the same spirit, we focus only on safety attacks with economic impact, namely conflicts among blocks that are accepted by clients [^1] Recall that the certificate which clients accept might differ with the certificate validators use to finalize.. Our goal is to ensure that, even if a client-side safety violation occurs, the system can recover and resume execution from a consistent state in which all honest clients harmed by the violation are economically reimbursed, without minting new coins [^2] Minting new coins inflates the currency and essentially penalizes honest validators..
Definition 0.
An SMR protocol regains economic restitution after safety violations with recovery parameter $\Delta_{R}$ if, after any safety violation, (i) all honest validators eventually recover to a common state $S$ for which there exists a verifiable state certificate that every honest validator witnesses at most $\Delta_{R}$ time after the equivocation, (ii) in state $S$, each client receives the full monetary value of every transaction it has accepted; (iii) validators responsible for the equivocation are punished, (iv) new coins are not minted.
2.3. SMR Game and Coalition-Compliance
To connect game-theoretic rational behavior to protocol-level resilience guarantees, we introduce the solution concept of $(n,k,f)$-coalition-compliance [^3] Coalition-compliance extends the notion of compliance originally introduced in (Karakostas et al., 2022).. Informally, a protocol is $(n,k,f)$-coalition-compliant if no rational validator prefers to follow a strategy that violates the SMR protocol, regardless of the actions of Byzantine validators. This property holds even if all rational and Byzantine validators collude.
By transforming an $(n,f)$-resilient protocol into an $(n,k,f)$-coalition-compliant protocol, we ensure that rational validators have no incentive to violate the SMR security properties. As a result, the transformed protocol effectively achieves $(n,k,f)$-resilience, even when rational and Byzantine parties collude.